OpenStudy (anonymous):

How many solutions does the following system of equations have? y=5/2x+2 2y=5x+8 One Two Zero Infinitely many

OpenStudy (anonymous):

i simplified the equations to y=5/2x+2 and y=5/2x+4

OpenStudy (anonymous):

same slope but different yint this meansss what ???